OpenStudy (anonymous):

please help me √a-7 -5=0 Step 1: Isolate the radical by adding 5 to both sides. Step 2: Square both sides and solve for a.

OpenStudy (anonymous):

Ok, so the first thing you are going to do is get rid of the 5. :) It doesn't belong with the "a" numbers and letters don't get along >:D So add it to each side.

OpenStudy (anonymous):

um is the -7 also under the radical?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh ok :)

OpenStudy (anonymous):

So when you add the 5 to both sides, you will end up with: \[\sqrt{a-7}=5\]

OpenStudy (anonymous):

Now that we have that, our next step is to get rid of the radical......now how do you think we do that?

OpenStudy (anonymous):

square both sides rt

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

So what would you get?

OpenStudy (anonymous):

thats the prob... we are getting a=32 so im not sure what im doing wrong

OpenStudy (anonymous):

Ok, so when you square it, you should get: a-7=25 then to get rid of the 7 you add it to both sides. Then you get: a=32 hmmmmm I'm pretty sure you're right?

OpenStudy (anonymous):

What are the answer choices?

OpenStudy (anonymous):

it doesnt have choices i just have to place the answer but it marked it wrong

OpenStudy (anonymous):

:( I'm pretty sure you're right..... i think you should email your teacher, maybe the computer system was wrong

OpenStudy (anonymous):

thinks so i have another one do you mine checking that one also

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.) \[\sqrt[4]{8z+9}=3\]

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i'm sorry, i'm not sure about how to solve this one :'( sorry

OpenStudy (anonymous):

k thank so much